OpenStudy (anonymous):

Use elimination to solve. 2x-3y=1 5x+4y=14 @cwrw238

OpenStudy (cwrw238):

multiply first equation by 5 and the seconnd by 2

OpenStudy (anonymous):

You mean make x = 5 and y = 2?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

no : 2x - 3y = 1 * 5 = 10x - 15y = 5 you multiply every term in the equation by 5 next you multiply second equation by 2

OpenStudy (anonymous):

Oh ok thanks!

OpenStudy (cwrw238):

the idea is to get 10x in each new equation so that subtracting these will eliminate x what do u get when multiply equation 2 by 2?

OpenStudy (anonymous):

you get 4.